OpenStudy (micahm):

State which set of three numbers could represent the lengths of the sides of a right triangle. A. 6,7,8 B. 16, 32 , 36 C. 5, 12 ,13 D. 8, 12, 16

OpenStudy (anonymous):

IDK... it is hard to say.

OpenStudy (micahm):

@ganeshie8

OpenStudy (micahm):

@mathmath333

OpenStudy (micahm):

@Compassionate

OpenStudy (anonymous):

i think it's A. 6,7,8

OpenStudy (micahm):

. Washington is 44 miles south of Lawrence. Midland is 63 miles east of Lawrence. How far apart are Washington and Midland? Round your answer to the nearest mile. A. 19 miles B. 27 MILES C. 77 MILES D. 107 MILES

OpenStudy (anonymous):

60 +40 =100 So I would say it's 107

OpenStudy (micahm):

What is the length of the diagonal of a rectangle whose width is 90 cm and whose length is 200 cm? (Round your answer to the nearest whole centimeter.) A. 214 CM B. 217 CM C. 219 CM D. 222 CM

OpenStudy (anonymous):

its not 107 its 77

OpenStudy (anonymous):

C. 77 miles

OpenStudy (micahm):

What is the length of the diagonal of a rectangle whose width is 90 cm and whose length is 200 cm? (Round your answer to the nearest whole centimeter.) A. 214 CM B. 217 CM C. 219 CM D. 222 CM

OpenStudy (mathmath333):

three numbers could represent the lengths of the sides of a right triangle C.) 5, 12 ,13 \(\large\tt \color{black}{\sqrt{5^2+12^2}}\) \(\large\tt \color{black}{=\sqrt{169}}\) \(\large\tt \color{black}{13}\)